# Interpolation theorems for conjugations and applications

### v1

@@ 2.1. Main results
(v)$\implies$(ii). Since the class of $C$-symmetric operators forms a subspace of ${\mathcal{B}}(H)$, we obtain, for every $k\in K$, $CN_{k}C=N_{k}^{*}$ and $C(x\otimes y)C=(x\otimes y)^{*}=y\otimes x$. Given that $x$ and $y$ have the same norm, the latter equality implies the existence of a unimodular $\alpha\in\mathbb{C}$ such that $Cx=\alpha y$.
Applying Theorem 2.3 (iv)$\implies$(ii) to $\{x_{i}\}_{i}=\{y_{i}\}_{i}=\{x\}$, we obtain a conjugation $J$ on $H$ such that $JN_{k}J=N_{k}^{*}$ for every $k\in K$ and $Jx=x$. Put $U=\overline{\alpha}CJ$; then, since $C$ and $J$ are both surjective isometries and $\alpha$ is unimodular, the linear operator $U:H\to H$ must be unitary. Moreover, we have
@@ 3. Interpolation theorems for conjugations : C-skew symmetric versions
A conjugate-linear operator $V$ on $H$ is said to be anti-unitary if it is a surjective isometry. Clearly, every conjugation on $H$ is anti-unitary.
@@
For $k\in\{1,2,\dots,\infty\}$, where $\infty=\aleph_{0}$, $H^{(k)}$ denotes the direct sum of $H$ with itself $k$ times. If $T\in{\mathcal{B}}(H)$, then $T^{(k)}\in\mathcal{B}(H^{(k)})$ is the direct sum of $T$ with itself $k$ times.
@@ 3.1. Main results
Let $N\in{\mathcal{B}}(H)$ be normal, and let $Z=X\cup Y$ where $X=\{x_{i}\}_{i\in I}$ and $Y=\{y_{i}\}_{i\in I}$ are orthogonal sets of non-zero vectors in $H$. Then, the following statements are equivalent:
@@
    There exists a conjugation $C$ on $H$ such that $CNC=-N^{*}$ and $Cx_{i}=y_{i}$ for every $i\in I$.
@@
    There exists an anti-unitary operator $V$ on $H$ such that $VNV^{-1}=-N^{*}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for every $i\in I$.
@@
    $N_{|L_{N,Z}^{\perp}}$ is skew-symmetric and, for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $i,j\in I$, we have
Before stating the second main result of this section, we recall some facts about normal operators on separable complex Hilbert spaces. Given a finite measure with compact support $\mu$ on $\mathbb{C}$, denote by $M_{\mu}$ the operator defined on $L^{2}(\mu)$ by $[M_{\mu}f](z)=zf(z)$ for all $f\in L^{2}(\mu)$. According to spectral multiplicity theory (see, for instance, [8, Page 298, Theorem 10.16]), each normal operator $N$ acting on $H$ is unitarily equivalent to
@@
In [17, Theorem 1.11], Li et al. proved that $N$ is skew-symmetric if and only if the sequence $\{\mu_{i}\}_{1\leq i\leq\infty}$ can be chosen so that, for every $i\in\{1,2,\dots,\infty\}$,
@@
Given a normal operator $N\in{\mathcal{B}}(H)$, we write $m(N)<\infty$ if the measure $\mu_{\infty}$ in (3.1) is zero; in other words, $M_{\mu_{\infty}}^{(\infty)}$ can be omitted from the decomposition (3.1). This definition makes sense due to the uniqueness property described in (3.2).
@@
Let $N\in{\mathcal{B}}(H)$ be a skew-symmetric normal operator, and let $Z$ be the set in Theorem 3.1 such that $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)<\infty$. Then, the following statements are equivalent:
@@
    There exists a conjugation $C$ on $H$ such that $CNC=-N^{*}$ and $Cx_{i}=y_{i}$ for every $i\in I$.
@@
In particular, if $I$ is finite, then $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)<\infty$ and the previous statements are equivalent.
@@
Suppose that $H=K_{1}\oplus K_{2}$, where $K_{1}$ and $K_{2}$ are infinite-dimensional separable complex Hilbert spaces, and let $\{e_{i}\}_{i\geq 1}$ and $\{f_{i}\}_{i\geq 1}$ be orthonormal bases of $K_{1}$ and $K_{2}$, respectively. For each $i\geq 1$, put $x_{i}=e_{i}\oplus 0$ and $y_{i}=0\oplus f_{i+1}$. Let $N$ be the normal operator on $H$ given by $N=I_{K_{1}}\oplus-I_{K_{2}}$. Clearly, $N\cong-N$; hence, by Lemma 3.5, $N$ is skew-symmetric.
For convenience, put $H_{1}=\vee\{x_{i}:i\geq 1\}$, $H_{2}=\vee\{y_{i}:i\geq 1\}$ and $H_{3}=\vee\{f_{1}\}$. Then, $H=H_{1}\oplus H_{2}\oplus H_{3}$ and $L_{N,Z}=H_{1}\oplus H_{2}$. Furthermore,
${N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}={N_{Z}}=\begin{bmatrix}I_{H_{1}}&0\\ 0&-I_{H_{2}}\end{bmatrix}\begin{array}[]{l}H_{1}\\ H_{2}\end{array}\cong M_{\delta_{1}}^{(\infty)}\oplus M_{\delta_{-1}}^{(\infty)}\cong\left(M_{\delta_{1}}\oplus M_{\delta_{-1}}\right)^{(\infty)}\cong M_{\mu}^{(\infty)}$
where $\mu=\delta_{1}+\delta_{-1}$ and $\delta_{a}$ denotes the Dirac measure, at $a\in\mathbb{C}$, on Borel subsets of $\mathbb{C}$. Thus, $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)\nless\infty$.
@@
    $E_{N}(\Delta)=I_{K_{1}}\oplus 0$ and $E_{N}(-\Delta)=0\oplus I_{K_{2}}$;
@@
    $E_{N}(\Delta)=0\oplus I_{K_{2}}$ and $E_{N}(-\Delta)=I_{K_{1}}\oplus 0$.
@@
Now, let us show that there exists no conjugation $C$ on $H$ satisfying the first assertion of Theorem 3.2. Suppose for contradiction that such a conjugation exists. Then,
$CH_{1}=C\vee\{x_{i}:i\geq 1\}=\vee\{Cx_{i}:i\geq 1\}=\vee\{y_{i}:i\geq 1\}=H_{2},$
and since $C^{2}=I_{H}$, we also have $CH_{2}=H_{1}$. As $C$ is a surjective isometry, it follows that $CH_{3}=C(H_{1}+H_{2})^{\perp}=(H_{1}+H_{2})^{\perp}=H_{3}$. Therefore, $C$ takes the form
$C=\begin{bmatrix}0&*&0\\ *&0&0\\ 0&0&C_{3}\end{bmatrix}\begin{array}[]{l}H_{1}\\ H_{2}\\ H_{3}\end{array}$
where $C_{3}$ is a conjugation on $H_{3}$. Finally, since $H_{3}$ reduces $N$ and $CNC=-N^{*}$, we obtain $-I_{H_{3}}=C_{3}(-I_{H_{3}})C_{3}=C_{3}N_{|H_{3}}C_{3}=-N_{|H_{3}}^{*}=I_{H_{3}}$; a contradiction.
When $X=\{x\}$, $Y=\{y\}$, and $\{x,y\}$ are linearly independent, we can add the following equivalent perturbation-type statement to Theorem 3.1.
@@
Let $N\in{\mathcal{B}}(H)$ be normal, and let $\{x,y\}$ be linearly independent vectors in $H$. The following statements are equivalent:
@@
    There exists a conjugation $C$ on $H$ satisfying $CNC=-N^{*}$ and $Cx=y$.
@@
    There exists a conjugation $C$ on $H$ such that $N+\lambda(x\otimes x-y\otimes y)$ is $C$-skew symmetric for every $\lambda\in\mathbb{C}$.
@@
(i)$\implies$(ii). Let $C$ be a conjugation satisfying (i). Then, for every $\lambda\in\mathbb{C}$, we have
$C\big{(}N+\lambda(x\otimes x-y\otimes y)\big{)}C$ $=CNC+\overline{\lambda}\big{(}(Cx)\otimes(Cx)-(Cy)\otimes(Cy)\big{)}$
$=-N^{*}+\overline{\lambda}(y\otimes y-x\otimes x)$
$=-\left(N+\lambda(x\otimes x-y\otimes y)\right)^{*}.$
@@
(ii)$\implies$(i). For convenience, put $T=x\otimes x-y\otimes y$. By the fact that the class of $C$-skew symmetric operators is a subspace of ${\mathcal{B}}(H)$, we clearly have $CNC=-N^{*}$ and $CTC=-T^{*}=-T$. Noting that $T$ is skew-symmetric, it must have a skew-symmetric matrix with respect to some orthonormal basis. In particular, since $T$ is a rank-two operator, its trace is well-defined and should be zero. Combining this with the fact that $T$ is self-adjoint, it follows that there exists orthonormal vectors $e_{1},e_{2}\in H$ and $\alpha\in\mathbb{R}$ such that
${\rm ran}(T)={\rm ker}(T-\alpha I)\oplus{\rm ker}(T+\alpha I),$
${\rm ker}(T-\alpha I)=\vee\{e_{1}\}\quad\mbox{and}\quad{\rm ker}(T+\alpha I)=\vee\{e_{2}\}.$
As $CTC=-T$, we get $C(T+\alpha I)C=-(T-\alpha I)$; hence
(3.4) $C(\vee\{e_{1}\})=C{\rm ker}(T-\alpha I)={\rm ker}(T+\alpha I)=\vee\{e_{2}\}.$
@@
Note that $x$ and $y$ have the same norm. In fact, $\langle x,x\rangle-\langle y,y\rangle$ is the trace of $T$. Hence, using [29, Theorem 2.1], we obtain a conjugation $J$ on $H$ satisfying $Jx=y$; one can check that $JTJ=-T$. The argument in (3.4) shows that $J(\vee\{e_{1}\})=\vee\{e_{2}\}$. Then, since $C$ and $J$ are isometries, there exist unimodular scalars $a,b\in\mathbb{C}$ such that $Ce_{1}=ae_{2}$ and $Je_{1}=be_{2}$. Furthermore, since $C^{2}=J^{2}=I$, we get $Ce_{2}=ae_{1}$ and $Je_{2}=be_{1}$. Noting that $C$ and $J$ are both conjugate-linear, we get $Ch=a\overline{b}Jh$ for every $h\in\vee\{e_{1},e_{2}\}$. In particular, since $x,y\in{\rm ran}(T)=\vee\{e_{1},e_{2}\}$, we obtain $Cx=a\overline{b}Jx=a\overline{b}y$. Put $\widetilde{C}=\overline{a}bC$; one can easily check that $\widetilde{C}$ is a conjugation that satisfies $\widetilde{C}N\widetilde{C}=-N^{*}$ and $\widetilde{C}x=y$. This ends the proof of the theorem. ∎
@@
3.2. Proof of Theorem 3.1 and 3.2
First, we begin with the following characterization of skew-symmetric normal operators.
@@
Let $N\in{\mathcal{B}}(H)$ be a normal operator. The following statements are equivalent:
@@
    There exists a unitary operator $U$ on $H$ such that $UNU^{*}=-N$.
@@
    There exists a conjugation $C$ on $H$ such that $CNC=-N^{*}$.
@@
    There exists an anti-unitary operator $V$ on $H$ such that $VNV^{-1}=-N^{*}$.
@@
The implication (ii)$\implies$(iii) is obvious.

### v3

@@ 2.1. Main results
(v)$\implies$(ii). Since the class of $C$-symmetric operators forms a subspace of ${\mathcal{B}}(\mathcal{H})$, we obtain, for every $k\in K$, $CN_{k}C=N_{k}^{*}$ and $C(x\otimes y)C=(x\otimes y)^{*}=y\otimes x$. Given that $x$ and $y$ have the same norm, the latter equality implies the existence of a unimodular $\alpha\in\mathbb{C}$ such that $Cx=\alpha y$.
Applying Theorem 2.4 (iii)$\implies$(i) to the particular case $\{x_{i}\}_{i}=\{y_{i}\}_{i}=\{x\}$, we obtain a conjugation $J$ on $\mathcal{H}$ such that $JN_{k}J=N_{k}^{*}$ for every $k\in K$ and $Jx=x$. Put $U=\overline{\alpha}CJ$; then, since $C$ and $J$ are both surjective isometries and $\alpha$ is unimodular, the linear operator $U:\mathcal{H}\to\mathcal{H}$ must be unitary. Moreover, we have
@@ 3. Interpolation theorems for conjugations : C-skew symmetric versions
A conjugate-linear operator $V$ on $\mathcal{H}$ is said to be anti-unitary if it is a surjective isometry. Clearly, every conjugation on $\mathcal{H}$ is anti-unitary.
@@
For $k\in\{1,2,\dots,\infty\}$, where $\infty=\aleph_{0}$, let $\mathcal{H}^{(k)}$ denote the direct sum of $\mathcal{H}$ with itself $k$ times. Accordingly, we denote by $T^{(k)}\in\mathcal{B}(\mathcal{H}^{(k)})$ the direct sum of $T\in{\mathcal{B}}(\mathcal{H})$.
@@ 3.1. Main results
The following statements are equivalent:
@@
    There exists a conjugation $C$ on $\mathcal{H}$ such that $CN_{k}C=-N_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
@@
    There exists an anti-unitary operator $V$ on $\mathcal{H}$ such that $VN_{k}V^{-1}=-N_{k}^{*}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for all $k\in K$ and $i\in I$.
@@
    There exists an anti-unitary operator $V_{1}$ on $\mathcal{L}_{1}:=\left(L_{\mathcal{N},X\cup Y}\right)^{\perp}$ such that
    $V_{1}{N_{k}}_{|\mathcal{L}_{1}}V_{1}^{-1}=-\left({N_{k}}_{|\mathcal{L}_{1}}\right)^{*}\quad\mbox{for every $k\in K$},$
    and for all finite subsets $K_{0}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, and $i,j\in I$ we have
    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},x_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(-\Delta_{k})\right)y_{j},y_{i}\right\rangle$
    and
    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},y_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(-\Delta_{k})\right)x_{j},y_{i}\right\rangle.$
Before stating the second main result of this section, we recall some facts about normal operators on separable complex Hilbert spaces. Given a finite measure with compact support $\mu$ on $\mathbb{C}$, denote by $M_{\mu}$ the operator defined on $L^{2}(\mu)$ by $[M_{\mu}f](z)=zf(z)$ for all $f\in L^{2}(\mu)$. According to spectral multiplicity theory (see, for instance, [9, p. 298, Theorem 10.16]), each normal operator $N$ acting on $\mathcal{H}$ is unitarily equivalent to
@@
In [18, Theorem 1.11], Li et al. proved that $N$ is skew-symmetric if and only if the sequence $\{\mu_{i}\}_{1\leq i\leq\infty}$ can be chosen so that, for every $i\in\{1,2,\dots,\infty\}$,
@@
Given a normal operator $N\in{\mathcal{B}}(\mathcal{H})$, we write $m(N)<\infty$ if the measure $\mu_{\infty}$ in (3.1) is zero; in other words, $M_{\mu_{\infty}}^{(\infty)}$ can be omitted from the decomposition (3.1). This definition makes sense due to the uniqueness property described in (3.2).
@@
Let $N\in{\mathcal{B}}(\mathcal{H})$ be a skew-symmetric normal operator, $N_{0}$ be the restriction of $N$ to $L_{\{N\},X\cup Y}$, and assume that $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$. The following statements are equivalent:
@@
    There exists a conjugation $C$ on $\mathcal{H}$ such that $CNC=-N^{*}$ and $Cx_{i}=y_{i}$ for every $i\in I$.
@@
In particular, if $I$ is finite, then $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$ and the previous statements are equivalent.
The proofs of Theorems 3.1 and 3.2 will be given at the end of this section.
@@
Suppose that $\mathcal{H}=\mathcal{K}_{1}\oplus\mathcal{K}_{2}$, where $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ are infinite-dimensional separable complex Hilbert spaces, and let $\{e_{i}\}_{i\geq 1}$ and $\{f_{i}\}_{i\geq 1}$ be orthonormal bases of $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$, respectively. For each $i\geq 1$, put $x_{i}=e_{i}\oplus 0$ $y_{i}=0\oplus f_{i+1}$. Let $N$ be the normal operator on $\mathcal{H}$ given by $N=I_{\mathcal{K}_{1}}\oplus-I_{\mathcal{K}_{2}}$. Clearly, $N\cong-N$; hence, by Lemma 3.7, $N$ is skew-symmetric.
For convenience, put $\mathcal{H}_{1}=\vee\{x_{i}:i\geq 1\}$, $\mathcal{H}_{2}=\vee\{y_{i}:i\geq 1\}$ and $\mathcal{H}_{3}=\vee\{0\oplus f_{1}\}$. Then, $\mathcal{H}=\mathcal{H}_{1}\oplus\mathcal{H}_{2}\oplus\mathcal{H}_{3}$ and $L_{\{N\},X\cup Y}=\mathcal{H}_{1}\oplus\mathcal{H}_{2}$. Furthermore,
${N_{0}}_{|{\rm ker}(N_{0})^{\perp}}={N_{0}}=\begin{bmatrix}I_{\mathcal{H}_{1}}&0\\ 0&-I_{\mathcal{H}_{2}}\end{bmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\end{array}\cong M_{\delta_{1}}^{(\infty)}\oplus M_{\delta_{-1}}^{(\infty)}\cong\left(M_{\delta_{1}}\oplus M_{\delta_{-1}}\right)^{(\infty)}\cong M_{\mu}^{(\infty)}$
where $\mu=\delta_{1}+\delta_{-1}$ and $\delta_{a}$ denotes the Dirac measure, at $a\in\mathbb{C}$, on Borel subsets of $\mathbb{C}$. Thus, $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)\nless\infty$.
@@
    $E_{N}(\Delta)=I_{\mathcal{K}_{1}}\oplus 0$ and $E_{N}(-\Delta)=0\oplus I_{\mathcal{K}_{2}}$;
@@
    $E_{N}(\Delta)=0\oplus I_{\mathcal{K}_{2}}$ and $E_{N}(-\Delta)=I_{\mathcal{K}_{1}}\oplus 0$.
@@
Now, let us show that there exists no conjugation $C$ on $\mathcal{H}$ satisfying the first assertion of Theorem 3.2. Suppose for contradiction that such a conjugation exists. Then,
$C\mathcal{H}_{1}=C\vee\{x_{i}:i\geq 1\}=\vee\{Cx_{i}:i\geq 1\}=\vee\{y_{i}:i\geq 1\}=\mathcal{H}_{2},$
and since $C^{-1}=C$, we also have $C\mathcal{H}_{2}=\mathcal{H}_{1}$, and hence $C(\mathcal{H}_{1}+\mathcal{H}_{2})=\mathcal{H}_{1}+\mathcal{H}_{2}$. As $C$ is a surjective isometry, it follows that $C\mathcal{H}_{3}=C(\mathcal{H}_{1}+\mathcal{H}_{2})^{\perp}=(\mathcal{H}_{1}+\mathcal{H}_{2})^{\perp}=\mathcal{H}_{3}$. Therefore, $C$ takes the form
$C=\begin{bmatrix}0&*&0\\ *&0&0\\ 0&0&C_{3}\end{bmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\\ \mathcal{H}_{3}\end{array}$
where $C_{3}$ is a conjugation on $\mathcal{H}_{3}$. Finally, since $\mathcal{H}_{3}$ reduces $N$ and $CNC=-N^{*}$, we obtain $-I_{\mathcal{H}_{3}}=C_{3}(-I_{\mathcal{H}_{3}})C_{3}=C_{3}N_{|\mathcal{H}_{3}}C_{3}=-N_{|\mathcal{H}_{3}}^{*}=I_{\mathcal{H}_{3}}$; a contradiction.
When $X=\{x\}$, $Y=\{y\}$, and $\{x,y\}$ are linearly independent, we can add the following statement to Theorem 3.1.
@@
Let $\{x,y\}$ be linearly independent vectors in $\mathcal{H}$. The following statements are equivalent:
@@
    There exists a conjugation $C$ on $\mathcal{H}$ satisfying $CN_{k}C=-N_{k}^{*}$, for every $k\in K$, and $Cx=y$.
@@
    There exists a conjugation $C$ on $\mathcal{H}$ such that $N_{k}+\lambda(x\otimes x-y\otimes y)$ is $C$-skew symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
@@
(i)$\implies$(ii). Let $C$ be a conjugation satisfying (i). Then, for all $k\in K$ and $\lambda\in\mathbb{C}$, we have
$C\big{(}N_{k}+\lambda(x\otimes x-y\otimes y)\big{)}C$ $=CN_{k}C+\overline{\lambda}\big{(}(Cx)\otimes(Cx)-(Cy)\otimes(Cy)\big{)}$
$=-N_{k}^{*}+\overline{\lambda}(y\otimes y-x\otimes x)$
$=-\left(N_{k}+\lambda(x\otimes x-y\otimes y)\right)^{*}.$
@@
(ii)$\implies$(i). For convenience, put $T=x\otimes x-y\otimes y$. By the fact that the class of $C$-skew symmetric operators is a subspace of ${\mathcal{B}}(\mathcal{H})$, we clearly have $CN_{k}C=-N_{k}^{*}$, for every $k\in K$, and $CTC=-T^{*}=-T$. Noting that $T$ is skew-symmetric, it must have a skew-symmetric matrix with respect to some orthonormal basis. In particular, since $T$ is a rank-two operator, its trace is well-defined and should be zero. Combining this with the fact that $T$ is self-adjoint, it follows that there exists orthonormal vectors $e_{1},e_{2}\in\mathcal{H}$ and $\alpha\in\mathbb{R}$ such that
${\rm ran}(T)={\rm ker}(T-\alpha I_{\mathcal{H}})\oplus{\rm ker}(T+\alpha I_{\mathcal{H}}),$
${\rm ker}(T-\alpha I_{\mathcal{H}})=\vee\{e_{1}\}\quad\mbox{and}\quad{\rm ker}(T+\alpha I_{\mathcal{H}})=\vee\{e_{2}\}.$
As $CTC=-T$, we get $C(T+\alpha I_{\mathcal{H}})C=-(T-\alpha I_{\mathcal{H}})$; hence
(3.4) $C(\vee\{e_{1}\})=C{\rm ker}(T-\alpha I_{\mathcal{H}})={\rm ker}(T+\alpha I_{\mathcal{H}})=\vee\{e_{2}\}.$
@@
Note that $x$ and $y$ have the same norm. In fact, $\langle x,x\rangle-\langle y,y\rangle$ is the trace of $T$. Hence, using [30, Theorem 2.1], we obtain a conjugation $J$ on $\mathcal{H}$ satisfying $Jx=y$; one can check that $JTJ=-T$. The argument in (3.4) shows that $J(\vee\{e_{1}\})=\vee\{e_{2}\}$. Then, since $C$ and $J$ are isometries, there exist unimodular scalars $a,b\in\mathbb{C}$ such that $Ce_{1}=ae_{2}$ and $Je_{1}=be_{2}$. Furthermore, since $C^{-1}=C$ and $J^{-1}=J$, we get $Ce_{2}=ae_{1}$ and $Je_{2}=be_{1}$. Noting that $C$ and $J$ are both conjugate-linear, we get $Ch=a\overline{b}Jh$ for every $h\in\vee\{e_{1},e_{2}\}$. In particular, since $x\in{\rm ran}(T)=\vee\{e_{1},e_{2}\}$, we obtain $Cx=a\overline{b}Jx=a\overline{b}y$. Put $\widetilde{C}=\overline{a}bC$; one can easily check that $\widetilde{C}$ is a conjugation that satisfies $\widetilde{C}N_{k}\widetilde{C}=-N_{k}^{*}$, for every $k\in K$, and $\widetilde{C}x=y$. This ends the proof of the theorem. ∎
Next, we shall apply the main results of this section to characterize skew symmetry of operators of the form
(3.5) $\sum_{j\in J}r_{j}(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})+{\rm i}B$
where the set $\{e_{j},f_{j}:j\in J\}\subset\mathcal{H}$ is orthonormal, $\{r_{j}\}_{j\in J}$ are distinct positive scalars and $B\in{\mathcal{B}}(\mathcal{H})$ is self-adjoint.
Before proceeding, it is worth noting that skew symmetric operators having the previous form are dense in ${\rm SSO}$. Let us prove this claim. Let $T\in{\mathcal{B}}(\mathcal{H})$ be a $C$-skew symmetric operator for some conjugation $C$ on $\mathcal{H}$. A quick examination of the cartesian decomposition of $T$ shows that we can write $T=A+{\rm i}B$ where $A$ and $B$ are $C$-skew symmetric self-adjoint operators. Using [5, Lemma 3.2], we obtain a sequence of self-adjoint diagonalizable $C$-skew symmetric operators $A_{n}$ that converges to $A$, with $\dim{\rm ker}(A_{n}-rI_{\mathcal{H}})=1$ whenever $r$ is an eigenvalue of $A_{n}$. Now, by skew-symmetry of $A_{n}$, we have $\dim{\rm ker}(A_{n}-rI_{\mathcal{H}})=\dim{\rm ker}(A_{n}+rI_{\mathcal{H}})$ for all real scalars $r$. Consequently, $T$ can be approximated by operators of the form (3.5).
Corollary 3.5.
Let $T\in{\mathcal{B}}(\mathcal{H})$ be an operator of the form (3.5) and denote by $B_{1}$ the restriction of $B$ to $\left(L_{\{B\},\{e_{j},f_{j}:j\in J\}}\right)^{\perp}$. Then, $T$ is skew-symmetric if and only if $B_{1}\cong-B_{1}$ and there exists unimodular scalars $\lambda_{j}$, for $j\in J$, such that
(3.6) $\lambda_{i}\left\langle E_{B}(\Delta)e_{i},e_{j}\right\rangle=\lambda_{j}\left\langle E_{B}(-\Delta)f_{j},f_{i}\right\rangle\;\mbox{and}\;\lambda_{i}\left\langle E_{B}(\Delta)e_{i},f_{j}\right\rangle=\lambda_{j}\left\langle E_{B}(-\Delta)e_{j},f_{i}\right\rangle$
for every Borel subset $\Delta$ of $\mathbb{C}$.
Proof.
Assume that (3.6) holds for every Borel subset $\Delta$ of $\mathbb{C}$. Since $B_{1}\cong-B_{1}$, Lemma 3.7 implies that $VB_{1}V^{-1}=-B_{1}^{*}$ for some anti-unitary operator $V$. Then, applying Theorem 3.1 to the self-adjoint operator $B$ and the orthonormal subsets $\{\lambda_{j}e_{j}\}_{j\in J}$ and $\{f_{j}\}_{j\in J}$, we obtain a conjugation $C$ on $\mathcal{H}$ such that $CBC=-B$ and $C(\lambda e_{j})=f_{j}$ for every $j\in J$. One can check that $C(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})C=-(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})=-(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})^{*}$ for every $j\in J$. Thus, $T$ is $C$-skew symmetric as the sum of two $C$-skew symmetric ones. This proves the sufficiency.
Suppose that $T$ is $C$-skew symmetric for some conjugation $C$ on $\mathcal{H}$, and denote $A=\sum_{j\in J}r_{j}(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})$. Then, $A$ and $B$ are both $C$-skew symmetric. Moreover, for every $j\in J$, ${\rm ker}(A-r_{j}I_{\mathcal{H}})=\vee\{e_{j}\}$ and ${\rm ker}(A+r_{j}I_{\mathcal{H}})=\vee\{f_{i}\}$. As in (3.4), we have that $Ce_{j}\in{\rm span}\{f_{j}\}$ for every $j\in J$. Hence, there exist unimodular scalars $\lambda_{j}$, for $j\in J$, so that $C(\lambda_{j}e_{j})=f_{j}$ for every $j\in J$. The necessity follows now by combining Theorem 3.1 and Lemma 3.7. ∎
Remark 3.6.
  1. (i)
    If the operator $T$ has rather the form
    $T=B+{\rm i}\sum_{j\in J}r_{j}(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})$
    with $\{e_{j},f_{j}\}_{j\in J}$, $\{r_{j}\}_{j\in J}$ and $B$ being as described in (3.5), then, taking into account that skew symmetry is invariant under scalar multiplication, the equivalence in the previous corollary still holds true.
  2. (ii)
    If $J$ is finite, then using Theorem 3.2, one can make minor changes in the proof of Corollary 3.5 to show that the condition “$B_{1}\cong-B_{1}$” can be replaced with the simpler one “$B\cong-B$”.
@@
3.2. Proof of Theorems 3.1 and 3.2
The following lemma provides conditions under which a family of mutually commuting normal operators have a common orthonormal basis in which their corresponding matrices are skew-symmetric. This result is inspired by [18, Theorem 1.10].
@@
Let $\{N_{k}\}_{k\in K}\subset{\mathcal{B}}(\mathcal{H})$ be mutually commuting normal operators. The following statements are equivalent:
@@
    There exists a conjugation $C$ on $\mathcal{H}$ such that
    $CN_{k}C=-N_{k}^{*}\quad\mbox{for every $k\in K$}.$
@@
    There exists an anti-unitary operator $V$ on $\mathcal{H}$ such that
    $VN_{k}V^{-1}=-N_{k}^{*}\quad\mbox{for every $k\in K$}.$
@@
    There exists a unitary operator $U$ on $\mathcal{H}$ such that
    $UN_{k}U^{*}=-N_{k}\quad\mbox{for every $k\in K$}.$
@@
The implication (i)$\implies$(ii) is obvious.